40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13180     Entry time: Wed Aug 9 19:21:18 2017
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS recovery 

Summary:

Between frequent MC1 excursions, I worked on ALS recovery today. Attachment #1 shows the out-of-loop ALS noise as of today evening (taken with arms locked to IR) - I have yet to check loop shapes of the ALS servos, looks like there is some tuning to be done.

On the PSL table:

  • First, I locked the arms to IR, ran the dither alignment servos to maximize transmission.
  • I used the IR beat PDs to make sure a beat existed, at approximately.
  • Then I used a scope to monitor the green beat, and tweaked steering mirror alignment until the beat amplitude was maximized. I was able to improve the X arm beat amplitude, which Koji and Naomi had tweaked last week, by ~factor of 2, and Y arm by ~factor of 10.
  • I used the DC outputs of the BBPDs to center the beam onto the PD.
  • Currently, the beat notes have amplitudes of ~-40dBm on the scopes in the control room (there are various couplers/amplifiers in the path so I am not sure what beatnote amplitude this translates to at the BBPD output). I have yet to do a thorough power budget, but I have in my mind that they used to be ~-30dBm. To be investigated.
  • Removed the fiber beat PD 1U chassis unit from the PSL table for further work. The fibers have been capped and remain on the PSL table. Cleaned the NW corner of the PSL table up a bit.

To do:

  • Optimization of the input pointing of the green beam for X (with PZTs) and Y (manual) arms.
  • ALS PDH servo loop measurement. Attachment #1 suggests some loop gain adjustment is required for both arms (although the hump centered around ~70Hz seem to be coming from the IR lock).
  • Power budgeting on the PSL table to compare to previous such efforts.

Note: Some of the ALS scripts are suffering from the recent inablilty of cdsutils to pull up testpoints (e.g. the script that is used to set the UGFs of the phase tracker servo). The workaround is to use DTT to open the test points first (just grab 0.1s time series for all channels of interest). Then the cdsutils scripts can read the required channels (but you have to keep the DTT open).

Attachment 1: ALS_oolSpec.pdf  72 kB  | Hide | Hide all
ALS_oolSpec.pdf
ELOG V3.1.3-